OpenStudy (andijo76):

Recently, two new transfer students joined your school, and both have received extremely high scores on the test. Which measure of central tendency would you use, and why? Explain which measure you would use in this situation by using actual numbers to illustrate your point. Calculate all three measures of central tendency and depict what happens before and after the addition of the two transfer students.

OpenStudy (timmyk14):

We make up numbers?

OpenStudy (andijo76):

i did the first half of the question but it seems like its asking the same question over

OpenStudy (timmyk14):

Can you show me the work you've done already?

OpenStudy (andijo76):

i will post the question first it had two parts this is the second

OpenStudy (andijo76):

You are part of a school and have to report the results of standardized test scores. Which measure of central tendency would you use? Assume that you don't know how your students will perform; so consider the general effects of outliers on the three measures of central tendency. Research and summarize the general effects that outliers have on each of the measures of central tendency. Then, using some real values, provide an example of how the three measures of central tendency vary due to outlier effects.

ganeshie8 (ganeshie8):

1) robust measures are : median, IQR 2) non robust measures are : mean, standard deviation

ganeshie8 (ganeshie8):

you would use robust measures when u have outliers in dataset

OpenStudy (andijo76):

mean= 90+90+95+95+98+92+94+100+100+97+96 =1145/12 =95.416666 =95.42 mean with outline= =90+90+95+95+98+98+92+94+100+100+97+96+100 =1245/12 =103.75 Median= (98+92)/12 =95 Mode= 95,100,90

ganeshie8 (ganeshie8):

you should choose LOW numbers so that the newly transferred students' scores get highlighted

ganeshie8 (ganeshie8):

maybe choose the scores such that the mean stays ~ 70 or so

ganeshie8 (ganeshie8):

and u can choose newly transferred students' marks around 95..

OpenStudy (andijo76):

so stay in the 70's and 80's

OpenStudy (nincompoop):

try 60's the styles then were good too